We are currently experiencing severe packet loss which is affecting connectivity to hosted web sites, email and servers.
Connectivity products such as broadband and leased lines are not affected.
Please accept our apologies for this outage and be assured that we are doing all we can to restore service as soon as possible.
Engineers are currently investigating. No ETR as yet. Please see comments below for updates.
One of our engineers in on his way to the data centre. ETA 20 minutes.
An engineer in the data centre is currently collaborating with our on-site engineer to determine the cause of the issue. No ETR is yet available.
Web sites and email services have ben restored. We are awaiting an update from the data centre engineers.
Engineers are still working on this issue.
Web sites and email services have ben restored. Engineers are still investigating the cause.
Engineers are still working on this issue. We will update as soon as possible.
The service was up and down throughout the morning and was permanently fixed at 2 pm. We again apologise for the outage which was caused by a DOS attack against the NHS.